# Aerial Active STAR-RIS-assisted Satellite-Terrestrial Covert Communications

### v1

@@ I Introduction
As a crucial component of the sixth-generation (6G) wireless communication systems, satellite-terrestrial communications have garnered significant attention in both military and civil sectors due to their extensive coverage and long-distance transmission capabilities [1]. For instance, in the aftermath of natural disasters such as earthquakes and floods, satellite-terrestrial systems can be an indispensable solution to re-establish communication links with disabled ground-based facilities. However, the persistent challenges of high path loss from the long-distance transmission and extreme penetration loss caused by obstacles continue to affect the performance of these systems, especially in dense urban environments [2].
@@
In such cases, low-altitude economy (LAE) offers new opportunities for the development of satellite-terrestrial communications [3]. As the core supporters of LAE, low-altitude platforms (LAPs), e.g, unmanned aerial vehicles (UAVs), can be deployed as the aerial relay nodes between satellites and ground users. Compared to the fixed relay communication mode, LAP-based aerial relay communications can bring an additional degree of freedom by dynamically adjusting their trajectories, thereby improving the channel capacity for the users in various locations. While the high line-of-sight (LoS) probability of air-to-ground channels improves the transmission performance, the inherent broadcast nature of wireless communications simultaneously heightens the security vulnerabilities [4]. As such, two widely used conventional techniques for ensuring secure communications are the encryption method in the high-layer network protocol stack and physical layer security [5]. However, these methods cannot fundamentally address security threats of satellite-terrestrial communications, particularly in the scenarios involving advanced technologies such as quantum computers, which could potentially break the conventional encryption methods, and more intelligent adversaries that may adapt to exploit emerging decoding techniques [6].
@@
Given the limitations of conventional security methods in addressing new forms of threats, covert communications have been recognized as a promising alternative [7]. Specifically, covert communications exploit the uncertainty of the Warden to prevent it from detecting the transmission by hiding the transmission information in noise, thereby providing a higher-level communication security. This approach aligns well with the dynamic capabilities of reconfigurable intelligent surfaces (RIS) and their variants, which adjust wireless channels by controlling intelligent reflection elements, thus positioning them as crucial technologies in covert communications for terrestrial environments [6]. Among the available RIS solutions, the active simultaneously transmitting and reflecting RIS (STAR-RIS) can become the preferred choice for mounting on LAPs in the satellite-terrestrial covert communication systems. This is because it can not only achieve the full-space coverage by a horizontal mounting mode, but also enhance the transmission rate under varying and complex channel conditions by actively amplifying and concentrating the wireless signal in the desired direction. Therefore, this paper aims to explore the satellite-terrestrial covert communications with the aid of the aerial active STAR-RIS (AASTAR-RIS).
@@ I-A1 Architecture of Satellite-Terrestrial Covert Communications
Existing works in the architecture of satellite-terrestrial covert communications can be mainly divided into direct mode and relay mode. For the direct mode, Zhang et al. [8] investigated an ultra-dense low Earth orbit (LEO) satellite downlink system, where the sidelobe leakages of other satellites are utilized to ensure the covertness. In [9], the authors developed a direct satellite-terrestrial downlink system by employing the rate-splitting multiple access technique to transmit jamming signals. However, this mode typically overlooks the long-distance path loss and complex propagation environments between satellites and ground users, which are particularly problematic in dense urban areas. In terms of the relay mode, Wu et al. [10] studied the ground device-based satellite-terrestrial covert relay communications, where multiple ground base stations are randomly selected to relay the wireless signal. Moreover, the authors in [11] presented a RIS-based satellite-terrestrial covert relay communication system, where a passive RIS is deployed on the surface of the building to optimize the signal reflection. However, the fixed position of relay platforms limits their adaptability in dynamic environments and reduces the flexibility of the system.
@@ I-A2 Uncertainty Source and Performance Metrics in Satellite-Terrestrial Covert Communications
In the field of satellite-terrestrial covert communications, several studies have explored modeling the uncertainty for the Warden from different sources and designing the corresponding performance metrics. For example, Guo et al. [12] presented a full-duplex ground receiver-based architecture of the satellite-terrestrial covert communications, where the transmit antenna of the ground receiver is used to send the jamming signal to raise the uncertainty of the Warden. In [13], the authors investigated the public message as a cover to produce the received power uncertainty in satellite-terrestrial covert communications, where the covert rate is maximized by jointly optimizing the overt message power allocation factor. Moreover, Feng et al. [14] proposed the passive RIS-assisted multi-satellite cooperative covert systems and utilized the active beamforming of multiple satellites and passive beamforming of RIS to achieve the communication covertness, while ensuring the maximization of the received signal power. However, these works typically focus on introducing uncertainty through additional components such as dedicated jamming devices, which overlook the inherent noise uncertainty of the environment itself. Moreover, the majority of existing works primarily center on optimizing the covert communication rate with limited consideration given to the fairness between users.
@@ I-A3 Optimization Methods for Satellite-Terrestrial Covert Communications
Existing studies on optimization methods for satellite-terrestrial covert communication systems primarily relied on convex optimization theory. For example, Hui et al. [15] proposed a two-stage iterative method with the block coordinate descent to maximize the average detection error probability (DEP). In [16], the authors utilized convex optimization to minimize the average age-of-information by taking the derivative of the transmission power of the transmitter in satellite-terrestrial covert communications. Moreover, Jia et al. [9] maximized the minimum covert rate by using the fractional programming, semidefinite relaxation and the S-procedure. However, these methods often assume static environments or require the channel state information (CSI) of the Warden. These assumptions limit the practicality of these methods in satellite-terrestrial covert communication systems, where the instantaneous CSI of the Warden is difficult to obtain and the detection threshold of the Warden may evolve over time.
@@
Reference (2022-2025) Satellite-Terrestrial Covert Communications Uncertainty Sources Performance Metrics Decision Variables Optimization Methods
Relay Mode RIS-assisted Relay Mode Movable RIS-assisted Relay Mode Environmental Noise Channel Capacity Fairness Trajectory Active Beamforming DRL Generative Models
[8]
[9]
[10]
[11]
[12]
[13]
[14]
[15]
[16]
@@ I-C Organization and Notations
[…]
@@ II-A Scenario Description
As shown in Fig. 1, we consider that a covert satellite-terrestrial downlink communication system, where a GEO satellite $a$ intends to communicate covertly with multiple ground users $\mathscr{K}\triangleq\{1,\dots,K\}$ against a passive ground Warden $w$. Due to the long distance and dense urban obstructions, we assume that the direct communication links between the GEO satellite and ground users are hindered by the low-gain receiving antennas and high propagation loss. To overcome this, the AASTAR-RIS [^1] The AASTAR-RIS can be horizontally mounted on various low aerial platforms, such as UAVs, electric vertical take-off and landing (eVTOL) aircraft or aerial vehicles, following the manner described in Ref. [17]. $r$ is deployed with $M=M_{x}\times M_{y}$ transmission and reflection elements, denoted by set $\mathscr{M}\triangleq\{1,\dots,M\}$, to enhance covert communication between the GEO satellite and ground users [^2] In some emergency scenarios, GEO satellites can offer robust and continuous communication coverage, making them a broadly applicable solution. Our future work will consider incorporating LEO satellites, which may be suited to certain specific scenarios..
Without loss of generality, three-dimensional Cartesian coordinates are adopted to denote the positions of all pertinent communication nodes. In our considered scenario, all of the GEO satellite, ground users, and Warden are assumed to be stationary, with their coordinates denoted as $\mathbf{q}_{a}=\left[x_{a},y_{a},z_{a}\right]^{\mathrm{T}}$, $\mathbf{q}_{k}=\left[x_{k},y_{k},0\right]^{\mathrm{T}}$, and $\mathbf{q}_{w}=\left[x_{w},y_{w},z_{w}\right]^{\mathrm{T}}$, respectively. Moreover, the total communication period $T$ is divided into $N$ time slots with an equal separation of $\delta_{t}=T/N$ and each time index is denoted as $n\in\mathscr{N}\triangleq\{1,\dots,N\}$. At time slot $n$, the position of the AASTAR-RIS is denoted as $\mathbf{q}_{r}[n]=\left[x_{r}[n],y_{r}[n],H\right]^{\mathrm{T}}$ with the fixed altitude $H$ to avoid collisions [7]. To facilitate analysis and optimization, $\delta_{t}$ is chosen to be sufficiently small, thereby ensuring that the network environment remains quasi-static within each time slot.
@@ II-B AASTAR-RIS Model
Due to the horizontal mounting configuration of the STAR-RIS, all elements of AASTAR-RIS are set in the full transmission mode. In this mode, the transmission coefficient matrix of the AASTAR-RIS at time slot $n$ can be given as follows [18]:
@@ II-C Channel Model
Let $\mathbf{h}_{ar}[n]$ and $\mathbf{h}_{rk}[n]$ stand for the channel gains from the GEO satellite to the AASTAR-RIS, from the AASTAR-RIS to the ground user $k$ and from the AASTAR-RIS to the Warden at time slot $n$, respectively. In this work, we consider that the channels from the GEO satellite to the AASTAR-RIS and from the AASTAR-RIS to the ground user $k$ and the Warden are Rician fading channels [19], which can be formulated at time slot $n$ as follows [20]:
@@
where $L_{0}$ represents the path loss at the reference distance of $1$ m, $d_{\mathcal{l}}[n]$ refers to the distance between the transmitter and receiver at time slot $n$, and $\kappa_{l}$ is the Rician factor. Moreover, $\mathbf{h}_{\mathcal{l}}^{\mathrm{LoS}}[n]$ and $\mathbf{h}_{\mathcal{l}}^{\mathrm{NLoS}}[n]$ are the LoS and the non-LoS (NLoS) components of the link $\mathcal{l}$, respectively. Specifically, the NLoS component can be modeled as the complex Gaussian distribution, i.e., $\mathbf{h}_{\mathcal{l}}^{\mathrm{NLoS}}[n]\sim\mathscr{CN}(\mathbf{0},\mathbf{I})$. For the communication link from the GEO satellite to the AASTAR-RIS, the LoS component at time slot $n$ can be expressed as follows [21]:
@@ II-E Binary Hypothesis Testing at Warden
Let the null hypothesis $\mathcal{H}_{0}$ and alternative hypothesis $\mathcal{H}_{1}$ denote the two circumstances that the GEO satellite is silent and sends the covert message to ground users, respectively. In particular, the Warden determines whether the GEO satellite is transmitting on the basis of the power of the signal received by a radiometer [6]. For the $i$-th channel at time slot $n$, the received signal at the Warden from the GEO satellite can be described as follows:
$\mathbf{y}_{w}^{i}[n]=\begin{cases}\begin{aligned} &\mathbf{h}_{rw}^{\mathrm{H}}[n]\mathbf{\Phi}[n]\left(\sqrt{p_{a}g_{a}}\mathbf{h}_{ar}[n]\mathbf{s}^{i}[n]+\mathbf{z}_{r}[n]\right)\\ &+z_{w}^{i}[n],\end{aligned}&\mathcal{H}_{1},\\ \begin{aligned} &z_{w}^{i}[n],\end{aligned}&\mathcal{H}_{0},\\ \end{cases}$ (8)
@@
In the binary testing problem, the decision rule of the Warden can be expressed as follows [20]:
@@
where $L$ represents the number of signal samples of the Warden, and $\tau[n]$ represents the decision threshold of the Warden at time slot $n$. Moreover, $\mathcal{D}_{0}$ and $\mathcal{D}_{1}$ are the decision results while supporting $\mathcal{H}_{0}$ and $\mathcal{H}_{1}$, respectively. Similar to [22], we consider $L\rightarrow\infty$, which means that an infinite amount of samples can be used at the Warden.
@@ III-A Covert Requirement
Due to variations of the wireless communication environment, the noise uncertainty of the Warden can be described as follows [22]:
$f_{\sigma_{w}^{2}}(x)=\begin{cases}\begin{aligned} &\frac{1}{2\ln(\rho)x}\end{aligned}&\text{if }\frac{1}{\rho}\hat{\sigma}_{w}^{2}\leq x\leq\rho\hat{\sigma}_{w}^{2},\\ \begin{aligned} &0,\end{aligned}&\text{otherwise},\\ \end{cases}$ (11)
@@
$\frac{\partial\xi[n]}{\partial\tau[n]}=\begin{cases}\begin{aligned} &-\frac{1}{(2\ln\rho)\tau[n]},\end{aligned}&\tau[n]\leq\iota[n]+\frac{\hat{\sigma}_{w}^{2}}{\rho},\\ \begin{aligned} &\frac{1}{2\ln\rho}\left(\frac{1}{\tau[n]-\iota[n]}-\frac{1}{\tau[n]}\right),\end{aligned}&\tau[n]>\iota[n]+\frac{\hat{\sigma}_{w}^{2}}{\rho}.\\ \end{cases}$ (13)
When $\tau[n]\leq\iota[n]+\frac{\hat{\sigma}{w}^{2}}{\rho}$, the derivative is negative, thereby implying that $\xi[n]$ decreases with respect to $\tau[n]$. When $\tau[n]>\iota[n]+\frac{\hat{\sigma}{w}^{2}}{\rho}$, the derivative is positive, and thus $\xi[n]$ increases with respect to $\tau[n]$. Moreover, the optimal detection threshold should be within $[\frac{\hat{\sigma}_{w}^{2}}{\rho},\rho\hat{\sigma}_{w}^{2}]$. Therefore, the optimal detection threshold for the Warden at time slot $n$ can be calculated as follows:
@@
$\xi^{*}[n]=\begin{cases}\begin{aligned} &1-\frac{1}{(2\ln\rho)x}\ln\left(1+\frac{\rho\iota[n]}{\hat{\sigma}_{w}^{2}}\right),\end{aligned}&\text{if }\iota[n]\leq\rho\hat{\sigma}_{w}^{2}-\frac{\hat{\sigma}_{w}^{2}}{\rho},\\ \begin{aligned} &0,\end{aligned}&\text{otherwise}.\\ \end{cases}$ (15)

### v2

@@ I Introduction
As a crucial component of the sixth-generation (6G) wireless communication systems, satellite-terrestrial communications have garnered significant attention in both military and civil sectors due to their extensive coverage and long-distance transmission capabilities [Mahboob2024]. For instance, in the aftermath of natural disasters such as earthquakes and floods, satellite-terrestrial systems can be an indispensable solution to re-establish communication links with disabled ground-based facilities. However, the persistent challenges of high path loss from the long-distance transmission and extreme penetration loss caused by obstacles continue to affect the performance of these systems, especially in dense urban environments [AlHourani2020].
@@
In such cases, low-altitude economy (LAE) offers new opportunities for the development of satellite-terrestrial communications [Jiahui2025]. As the core supporters of LAE, low-altitude platforms (LAPs), e.g, unmanned aerial vehicles (UAVs), can be deployed as the aerial relay nodes between satellites and ground users. Compared to the fixed relay communication mode, LAP-based aerial relay communications can bring an additional degree of freedom by dynamically adjusting their trajectories, thereby improving the channel capacity for the users in various locations. While the high line-of-sight (LoS) probability of air-to-ground channels improves the transmission performance, the inherent broadcast nature of wireless communications simultaneously heightens the security vulnerabilities [Ding2024]. As such, two widely used conventional techniques for ensuring secure communications are the encryption method in the high-layer network protocol stack and physical layer security [li2024two]. However, these methods cannot fundamentally address security threats of satellite-terrestrial communications, particularly in the scenarios involving advanced technologies such as quantum computers, which could potentially break the conventional encryption methods, and more intelligent adversaries that may adapt to exploit emerging decoding techniques [Chen2023].
@@
Given the limitations of conventional security methods in addressing new forms of threats, covert communications have been recognized as a promising alternative [Wang2023]. Specifically, covert communications exploit the uncertainty of the Warden to prevent it from detecting the transmission by hiding the transmission information in noise, thereby providing a higher-level communication security. This approach aligns well with the dynamic capabilities of reconfigurable intelligent surfaces (RIS) and their variants, which adjust wireless channels by controlling intelligent reflection elements, thus positioning them as crucial technologies in covert communications for terrestrial environments [Chen2023]. Among the available RIS solutions, the active simultaneously transmitting and reflecting RIS (STAR-RIS) can become the preferred choice for mounting on LAPs in the satellite-terrestrial covert communication systems. This is because it can not only achieve the full-space coverage by a horizontal mounting mode, but also enhance the transmission rate under varying and complex channel conditions by actively amplifying and concentrating the wireless signal in the desired direction. Therefore, this paper aims to explore the satellite-terrestrial covert communications with the aid of the aerial active STAR-RIS (AASTAR-RIS).
@@ I-A1 Architecture of Satellite-Terrestrial Covert Communications
Existing works in the architecture of satellite-terrestrial covert communications can be mainly divided into direct mode and relay mode. For the direct mode, Zhang et al. [Zhanglei2024] investigated an ultra-dense low Earth orbit (LEO) satellite downlink system, where the sidelobe leakages of other satellites are utilized to ensure the covertness. In [Jia2025], the authors developed a direct satellite-terrestrial downlink system by employing the rate-splitting multiple access technique to transmit jamming signals. However, this mode typically overlooks the long-distance path loss and complex propagation environments between satellites and ground users, which are particularly problematic in dense urban areas. In terms of the relay mode, Wu et al. [Wu2022] studied the ground device-based satellite-terrestrial covert relay communications, where multiple ground base stations are randomly selected to relay the wireless signal. Moreover, the authors in [Song2023] presented a RIS-based satellite-terrestrial covert relay communication system, where a passive RIS is deployed on the surface of the building to optimize the signal reflection. However, the fixed position of relay platforms limits their adaptability in dynamic environments and reduces the flexibility of the system.
@@ I-A2 Uncertainty Source and Performance Metrics in Satellite-Terrestrial Covert Communications
In the field of satellite-terrestrial covert communications, several studies have explored modeling the uncertainty for the Warden from different sources and designing the corresponding performance metrics. For example, Guo et al. [Guo2024] presented a full-duplex ground receiver-based architecture of the satellite-terrestrial covert communications, where the transmit antenna of the ground receiver is used to send the jamming signal to raise the uncertainty of the Warden. In [Yu2024], the authors investigated the public message as a cover to produce the received power uncertainty in satellite-terrestrial covert communications, where the covert rate is maximized by jointly optimizing the overt message power allocation factor. Moreover, Feng et al. [Feng2024] proposed the passive RIS-assisted multi-satellite cooperative covert systems and utilized the active beamforming of multiple satellites and passive beamforming of RIS to achieve the communication covertness, while ensuring the maximization of the received signal power. However, these works typically focus on introducing uncertainty through additional components such as dedicated jamming devices, which overlook the inherent noise uncertainty of the environment itself. Moreover, the majority of existing works primarily center on optimizing the covert communication rate with limited consideration given to the fairness between users.
@@ I-A3 Optimization Methods for Satellite-Terrestrial Covert Communications
Existing studies on optimization methods for satellite-terrestrial covert communication systems primarily relied on convex optimization theory. For example, Hui et al. [Hui2024] proposed a two-stage iterative method with the block coordinate descent to maximize the average detection error probability (DEP). In [Cao2024], the authors utilized convex optimization to minimize the average age-of-information by taking the derivative of the transmission power of the transmitter in satellite-terrestrial covert communications. Moreover, Jia et al. [Jia2025] maximized the minimum covert rate by using the fractional programming, semidefinite relaxation and the S-procedure. However, these methods often assume static environments or require the channel state information (CSI) of the Warden. These assumptions limit the practicality of these methods in satellite-terrestrial covert communication systems, where the instantaneous CSI of the Warden is difficult to obtain and the detection threshold of the Warden may evolve over time.
@@
Reference
(2022-2025)
Satellite-Terrestrial
Covert Communications
Uncertainty
Sources
Performance Metrics Decision Variables Optimization Methods
Relay
Mode
RIS-assisted
Relay Mode
Movable RIS-assisted
Relay Mode
Environmental
Noise
Channel
Capacity
Fairness Trajectory
Active
Beamforming
DRL
Generative
Models
[Zhanglei2024]
[Jia2025]
[Wu2022]
[Song2023]
[Guo2024]
[Yu2024]
[Feng2024]
[Hui2024]
[Cao2024]
@@ I-C Organization and Notations
Notations: We use plain, boldface lowercase, boldface uppercase, and calligraphic symbols to represent scalars, vectors, matrices, and sets (e.g., $a$, $\mathbf{a}$, $\mathbf{A}$, and $\mathcal{A}$), respectively. For a vector or matrix, the Euclidean and Frobenius norm are represented as $\|\cdot\|$ and the transpose and conjugate transpose are denoted as $\mathbf{\{\cdot\}}^{\mathrm{T}}$ and $\mathbf{\{\cdot\}}^{\mathrm{H}}$, respectively. $\mathrm{diag}(\mathbf{a})$ defines a diagonal matrix whose diagonal elements originate from $\mathbf{a}$. We denote the absolute value of a complex variable by $|\cdot|$, the Kronecker product by $\otimes$, and the identity matrix by $\mathbf{I}$. In addition, we denote the real-value and circularly symmetric complex Gaussian distribution with mean $\bm{\mu}$ and covariance matrix $\bm{\Lambda}$ by $\mathscr{CN}(\bm{\mu},\bm{\Lambda})$ and $\mathscr{N}(\bm{\mu},\bm{\Lambda})$, respectively. In addition, $\Pr(\cdot)$ represents the probability of an event, and $\mathbb{E}[\cdot]$ denotes the expected value of a random variable.
@@ II-A Scenario Description
As shown in Fig. 1, we consider that a covert satellite-terrestrial downlink communication system, where a GEO satellite $a$ intends to communicate covertly with multiple ground users $\mathscr{K}\triangleq\{1,\dots,K\}$ against a passive ground Warden $w$. Due to the long distance and dense urban obstructions, we assume that the direct communication links between the GEO satellite and ground users are hindered by the low-gain receiving antennas and high propagation loss. To overcome this, the AASTAR-RIS [^1] The AASTAR-RIS can be horizontally mounted on various low aerial platforms to ensure wide elevation-angle coverage and reduce aerodynamic drag during horizontal mobility, such as UAVs, electric vertical take-off and landing (eVTOL) aircraft or aerial vehicles, following the manner described in Ref. [WDLukito2024]. $r$ is deployed with $M=M_{x}\times M_{y}$ transmission and reflection elements, denoted by set $\mathscr{M}\triangleq\{1,\dots,M\}$, to enhance covert communication between the GEO satellite and ground users [^2] In some emergency scenarios, GEO satellites can offer robust and continuous communication coverage, making them a broadly applicable solution. Our future work will consider incorporating LEO satellites, which may be suited to certain specific scenarios..
Without loss of generality, three-dimensional Cartesian coordinates are adopted to denote the positions of all pertinent communication nodes. In our considered scenario, all of the GEO satellite, ground users, and Warden are assumed to be stationary, with their coordinates denoted as $\mathbf{q}_{a}=\left[x_{a},y_{a},z_{a}\right]^{\mathrm{T}}$, $\mathbf{q}_{k}=\left[x_{k},y_{k},0\right]^{\mathrm{T}}$, and $\mathbf{q}_{w}=\left[x_{w},y_{w},z_{w}\right]^{\mathrm{T}}$, respectively. Moreover, the total communication period $T$ is divided into $N$ time slots with an equal separation of $\delta_{t}=T/N$ and each time index is denoted as $n\in\mathscr{N}\triangleq\{1,\dots,N\}$. At time slot $n$, the position of the AASTAR-RIS is denoted as $\mathbf{q}_{r}[n]=\left[x_{r}[n],y_{r}[n],H\right]^{\mathrm{T}}$ with the fixed altitude $H$ to avoid collisions [Wang2023]. To facilitate analysis and optimization, $\delta_{t}$ is chosen to be sufficiently small, thereby ensuring that the network environment remains quasi-static within each time slot. Moreover, for our considered scenario, we make the following assumptions about CSI as follows.
  • CSI available to the AASTAR-RIS: The AASTAR-RIS is assumed to have access to the instantaneous CSI from the GEO satellite and all ground users, which can be accomplished by well-established techniques in cooperative wireless communication systems [Zheng2022, Guo2025]. However, since we model the Warden as a passive and non-cooperative adversary, the CSI from the AASTAR-RIS to the Warden is assumed to be entirely unknown to the AASTAR-RIS.
  • CSI available to the Warden: To ensure rigorous covertness constraint modeling, we assume that the Warden possesses perfect instantaneous CSI for all relevant links, as adopted in [Shahzad2018, He2025].
@@ II-B AASTAR-RIS Model
Due to the horizontal mounting configuration of the STAR-RIS, all elements of AASTAR-RIS are set in the full transmission mode. In this mode, the transmission coefficient matrix of the AASTAR-RIS at time slot $n$ can be given as follows [Mu2022]:
@@ II-C Channel Model
Let $\mathbf{h}_{ar}[n]$ and $\mathbf{h}_{rk}[n]$ stand for the channel gains from the GEO satellite to the AASTAR-RIS, from the AASTAR-RIS to the ground user $k$ and from the AASTAR-RIS to the Warden at time slot $n$, respectively. In this work, we consider that the channels from the GEO satellite to the AASTAR-RIS and from the AASTAR-RIS to the ground user $k$ and the Warden are Rician fading channels [Zheng2025], which can be formulated at time slot $n$ as follows [Wang2025]:
@@
where $L_{0}$ represents the path loss at the reference distance of $1$ m, $d_{\mathcal{l}}[n]$ refers to the distance between the transmitter and receiver at time slot $n$, and $\kappa_{l}$ is the Rician factor. Moreover, $\mathbf{h}_{\mathcal{l}}^{\mathrm{LoS}}[n]$ and $\mathbf{h}_{\mathcal{l}}^{\mathrm{NLoS}}[n]$ are the LoS and the non-LoS (NLoS) components of the link $\mathcal{l}$, respectively. Specifically, the NLoS component can be modeled as the complex Gaussian distribution, i.e., $\mathbf{h}_{\mathcal{l}}^{\mathrm{NLoS}}[n]\sim\mathscr{CN}(\mathbf{0},\mathbf{I})$. For the communication link from the GEO satellite to the AASTAR-RIS, the LoS component at time slot $n$ can be expressed as follows [Lu2021]:
@@ II-E Binary Hypothesis Testing at Warden
Let the null hypothesis $\mathcal{H}_{0}$ and alternative hypothesis $\mathcal{H}_{1}$ denote the two circumstances that the GEO satellite is silent and sends the covert message to ground users, respectively. In particular, the Warden determines whether the GEO satellite is transmitting on the basis of the power of the signal received by a radiometer [Chen2023]. For the $i$-th channel at time slot $n$, the received signal at the Warden from the GEO satellite can be described as follows:
[…]
@@
In the binary testing problem, the decision rule of the Warden can be expressed as follows [Wang2025]:
@@
where $L$ represents the number of signal samples of the Warden, and $\tau[n]$ represents the decision threshold of the Warden at time slot $n$. Moreover, $\mathcal{D}_{0}$ and $\mathcal{D}_{1}$ are the decision results while supporting $\mathcal{H}_{0}$ and $\mathcal{H}_{1}$, respectively. Similar to [Zhang2024], we consider $L\rightarrow\infty$, which means that an infinite amount of samples can be used at the Warden.
@@ III-A Covert Requirement
Due to variations of the wireless communication environment, the noise uncertainty of the Warden can be described as follows [Zhang2024]:
[…]
@@
When $\tau[n]\leq\iota[n]+\frac{\hat{\sigma}{w}^{2}}{\rho}$, the derivative is negative, thereby implying that $\xi[n]$ decreases with respect to $\tau[n]$. When $\tau[n]>\iota[n]+\frac{\hat{\sigma}{w}^{2}}{\rho}$, the derivative is positive, and thus $\xi[n]$ increases with respect to $\tau[n]$. Moreover, the optimal detection threshold should be within $\frac{\hat{\sigma}_{w}^{2}}{\rho},\rho\hat{\sigma}_{w}^{2}]$. Therefore, the optimal detection threshold for the Warden at time slot $n$ can be calculated as follows:
@@
[…]